OpenStudy (anonymous):

1. Solve 3x - x + 2 = 12. A. -5 B. 4 C. 7 D. 5 2. Solve 2(x + 1) + 4 = 8. A. 1 B. 6.5 C. 4 D. 3 3. Solve for Y/-4 + 5 = 15 A. 80 B. -80 C. 40 D. -40

OpenStudy (anonymous):

2x+2=12 2x=12-2 2x=10 x=5

OpenStudy (anonymous):

3x - x + 2 = 12. i.e. 2x=12-2 i.e. 2x= 10 i.e. x=10/2 i.e. x=5

OpenStudy (anonymous):

2x+2+4=8 2x+6=8 2x=8-6 2x=2 x=1

OpenStudy (anonymous):

Im not really sure about number 3 i suck at fractions.

OpenStudy (anonymous):

 is dis the question first tell me